Grade 12Algebra

Give the steps you will use to separate the variable and then solve the equation:
30p/10=6

Pawan Prajapati
5 Years agoGrade 12

2 Answers

Harshit Singh
5 Years ago
Dear Student

First we have to multiply both sides of the equation by 10,
Then, we get,
= (3p/10) × 10 = 6 × 10
= 3p = 60
Now,
We have to divide both sides of the equation by 3,
Then, we get,
= 3p/3 = 60/3
= p = 20

Thanks
Prachi Singh
5 Years ago
Dear Student,
 
We can solve the following expression as follows :
30p/10 = 6
30p = 6×10
30p = 60
P = 60/30
P = 2
 
Hope it helps!
Thank you